OpenStudy (anonymous):

i have answered the question of the time to reach the greatest voltage rate of change with the equation v=100sin (200pit+pi/4). My answer is t=-0.00125. However i have to evaluate t from the following expression, d2v/dt2 = - (4x10^6 pi/2)sin (200 pit + pi/4) = 0. How do i do this?

OpenStudy (rational):

take the integral

OpenStudy (rational):

\(\large \frac{d^2v}{dt^2} = -(4*10^6\pi/2) \sin (200 \pi t + \pi/4)\)

OpenStudy (rational):

\(\large \frac{dv}{dt} = \int -(4*10^6\pi/2) \sin (200 \pi t + \pi/4) dt\)

OpenStudy (anonymous):

i dont understand how i re-write it because my next part was dv/dt is max when cos(200pit + pi/4) is 1 or -1 so t =-1/800 ot 1/800 which gives t= -0.00125 or 0.00125

ganeshie8 (ganeshie8):

yes, take the integral first

ganeshie8 (ganeshie8):

\(\large \frac{dv}{dt} = \int -(4*10^6\pi/2) \sin (200 \pi t + \pi/4) dt\) \(\large ~~~~ = -(4*10^6\pi/2)\int \sin (200 \pi t + \pi/4) dt\) \(\large ~~~~ = [4*10^6\pi/(2*200\pi)] \cos (200 \pi t + \pi/4) \) \(\large ~~~~ = 10^4 \cos (200 \pi t + \pi/4) \)

ganeshie8 (ganeshie8):

this reaches max-value, when cos term becomes 1

ganeshie8 (ganeshie8):

so u need to solve : \(\cos (200 \pi t + \pi/4) = 1\) for \(t\)

ganeshie8 (ganeshie8):

if that makes any sense...

OpenStudy (anonymous):

struggling with this sorry, how do i make cos (200pit+pi/4) show for t

ganeshie8 (ganeshie8):

take \(\cos^{-1}\) both sides

ganeshie8 (ganeshie8):

\(\cos (200 \pi t + \pi/4) = 1\) take \(\cos^{-1}\) both sides : \(\cos^{-1}\cos (200 \pi t + \pi/4) = \cos^{-1}(1)\) \(200 \pi t + \pi/4 = 0\) solve \(t\)

OpenStudy (anonymous):

so is the answer t = -0.00125 or 0.00125seconds?

OpenStudy (anonymous):

thanks for all your help

ganeshie8 (ganeshie8):

yes, becoz cos(x) = cos(-x) u need to take both positive and negative values

ganeshie8 (ganeshie8):

u wlc :)